OpenStudy (anonymous):

Which equation best represents the regression line for the data given in the table above? y = –3x – 20 y = –3x + 20 y = 3x + 20 y = 3x – 20

OpenStudy (anonymous):

x y 1 13 2 17 3 14 4 8 5 3 6 –2 7 2

OpenStudy (anonymous):

put the x value into the equation, we can get the value of y, and see which one is close to the y from the list

OpenStudy (anonymous):

y = –3(1) – 20 y = -23 y = –3(1) + 20 y = 17 y = 3(1) + 20 y = 23 y = 3(1) – 20 y = -17

OpenStudy (anonymous):

good, so which one is closest to 13?

OpenStudy (anonymous):

B y = –3(1) + 20 y = 17

OpenStudy (anonymous):

:) you got it! but remember, usually you need put at least three x values into the equation to make sure this is the right answer